OpenStudy (anonymous):

prove that limx→∞ (arctanx/x ) = ∞. solve this one prove that limx→∞ (arctanx/x ) = ∞. solve this one @Mathematics

OpenStudy (jamesj):

The limit is not infinity, but zero.

OpenStudy (perl):

lim arctan(00) / 00 = lim pi/4 / 00 = 0

OpenStudy (anonymous):

the quation is to prove,,,,

OpenStudy (anonymous):

the limit is infinite,,,

OpenStudy (perl):

no it isnt

OpenStudy (perl):

prove? do you have a 100 dollars ? lol

OpenStudy (perl):

hassank, use properties of limits

OpenStudy (anonymous):

sure

OpenStudy (across):

\[\lim_{x\to\infty}\frac{\arctan(x)}{x}=\frac{\lim_{x\to\infty}\arctan(x)}{\lim_{x\to\infty}x}=\frac{\pi/2}{\infty}=0.\]

OpenStudy (perl):

yes she is right

OpenStudy (anonymous):

she is right,,,but iam right toooo

OpenStudy (perl):

you can prove seperately lim arctan x goes to pi/4

OpenStudy (perl):

errr, pi/2 ?

OpenStudy (anonymous):

the lmit can be infinite to

OpenStudy (jamesj):

\[ \lim_{x \rightarrow \infty} \arctan(x) = \pi/2 \] and \[ \lim_{x \rightarrow \infty} 1/x = 0 \] Therefore \[\lim_{x \rightarrow \infty} \arctan(x) /x = \lim_{x \rightarrow \infty} \arctan(x) \ \ . \ \ \lim_{x \rightarrow \infty} 1/x = (\pi/2) \ .\ 0 = 0 \]

OpenStudy (anonymous):

but my quation is show that the limit can be infite,,,i know your answer is correct,,,but show that it can be infinite

OpenStudy (across):

I know that 2 + 2 = 4, but show that 2 + 2 = 100000000 as well.

OpenStudy (anonymous):

what across said

OpenStudy (anonymous):

i have got this quation from my teacher,,,

OpenStudy (perl):

if x->0 , i believe it is going to infinity

OpenStudy (anonymous):

f(x) = arctanx/x ; g(x) = 1/x make a graph of both functions,,,,and solve the joining point which is tan x after that you se the graph of f(x) > g(x) after the joining point,,, then make integral from tanx to infinite ((f(x)-g(x)) and you will that the limit ends infinite

OpenStudy (perl):

and?